# Robust Denoising Neural Reranker for Recommender Systems

### v1

@@ Abstract
For multi-stage recommenders in industry, a user request would first trigger a simple and efficient retriever module that selects and ranks a list of relevant items, then calls a slower but more sophisticated deep reranking model that refines the item arrangement before exposure to the user. The latter model typically reranks the item list conditioned on the users’ history content and the initial ranking from retrievers. Although this two-stage retrieval-ranking framework demonstrates practical effectiveness, the significance of retriever scores from the previous stage has been limitedly explored, which is informative. In this work, we first theoretically analyze the limitations of using retriever scores as the reranker’s input directly and argue that the reranking task is essentially a noise reduction problem from the retriever scores. Following this notion, we derive an adversarial framework, DNR, that associates the denoising reranker with a carefully designed noise generation module. We extend the conventional score error minimization term with three augmented objectives, including: 1) a denoising objective that aims to denoise the noisy retriever scores to align with the user feedback; 2) an adversarial retriever score generation objective that improves the exploration in the retriever score space; and 3) a distribution regularization term that aims to align the distribution of generated noisy retriever scores with the real ones. Extensive experiments are conducted on three public datasets, together with analytical support, validating the effectiveness of the proposed DNR.
@@ 1 Introduction
Recommender systems aim to personalize a list of items that maximizes user engagement signals such as purchases, video plays, and link sharing. To achieve this, industrial solutions [1, 2, 3] typically employ a multi-stage architecture comprising a retriever and a reranker. The retriever, which may further decompose into several consecutive sub-stages of retrieval and ranking models, efficiently filters a large item pool down to a manageable candidate set of relevant items. Then the reranker adopts a more sophisticated but more accurate model that refines the candidate item set into an optimal ordering for the list-wise exposure. Formally, the goal of the latter reranking task [4, 5, 6, 7] is to learn an ordering function over the retrieved candidates that maximizes user responses on the top-reranked items. This task is particularly challenging due to the combinatorial nature of the item permutation space. Exhaustively searching for the optimal list exposure in this space becomes infeasible as the candidate set grows, making this a key bottleneck in real-world recommender systems.
@@
To overcome this challenge, early studies propose to refine the initial retriever scores [5, 8, 6] by modeling the item mutual influences in the candidate item list. Recent state-of-the-art solutions find it more effective to frame the reranking task as a generative modeling task [9, 10, 11] where the list generation probability should directly align with the user’s list-wise rewards. Yet, these methods largely overlook the potential of initial retriever scores, which may offer rich prior information to enhance the reranking performance. In practice, a straightforward approach to leverage initial retriever scores is to include them as input features directly, which shows promise to improve reranking performance as in Figure 1(a).
Reranking as a noise reduction problem: Though effective, this naive solution neglects the noisy nature of the initial retriever scores, which may cause model robustness concerns and downgrade the effectiveness of the learned reranking model. Specifically, the retriever stage, constrained by strict latency and computational budgets, usually adopts a simpler model than the reranker. In contrast, the reranker, operating on a much smaller candidate set, can use a more sophisticated architecture and thus achieves significantly higher accuracy (Figure 1(b)). In other words, by assuming higher noise in the retriever scores, it naturally indicates a noise reduction process of the reranking stage. This noise reduction pattern is more evident when we observe the changes in the embedding representations of ground truth target items, comparing Figure 1(c) and Figure 1(d). This means that the reranker behaves like a denoising model even without explicit modeling of the retriever’s noise. However, as we will show in section 3.1, the retriever’s noise theoretically contributes to the misalignment between the traditional objective of reranking and the user response likelihood maximization goal, restricting the generalization ability of the reranker. As a countermeasure, we need a more robust solution that can learn a denoising reranker that is aware of the retriever’s noisy score distribution, while taking advantage of the information from retriever scores.
@@
The proposed solution: Based on the previous analysis, we propose to augment the denoising neural reranker (DNR) with a carefully designed noise injection model, which estimates the initial retriever score’s posterior distribution. We theoretically show that the data likelihood maximization goal derives a solution framework that includes three learning objectives: 1) an augmented denoise objective that aims to boost the user feedback reconstruction under unseen retriever scores, increasing the reranker’s effectiveness to denoise noisy retriever scores; 2) an adversarial retriever score generation objective that aims to generate irregular retriever scores for the reranker that are hard to denoise, which in turn improves the reranker’s robustness; and 3) a score posterior regularization term that aims to align the noise-aware posterior model with the initial retriever score distribution, guaranteeing data alignment with the system behavior in previous stages. Additionally, we show that the noise-injecting posterior and the denoising reranker collaborate with each other in a similar way to the one-step diffusion framework [12], but standard Gaussian noise is not necessarily an empirically optimal choice. Instead, we illustrate that the binomial nature of the user response better suits a Beta distribution in the retriever score space, which derives our heuristic solution for the posterior. Still, this naive design lacks personalization abilities. To mitigate this limitation, we design a learnable posterior model that approximately optimizes the three augmented learning objectives.
Empirical supports and findings: We conduct empirical experiments on three public datasets to verify the superiority of DNR against leading rerankers. Notably, the reranker paired with the learnable posterior best aligns with the observed ground-truth posterior, which theoretically improves the alignment between the reranker’s learning objective and the user feedback data likelihood maximization goal, resulting in the best empirical reranking performance in our experiment.
In general, this work provides evidence that the initial retriever scores may be more informative than we think, and a more comprehensive utilization of this information may derive a solution that surpasses state-of-the-art methods.
@@ 2.1 Multi-stage Recommendation
Modern content-intensive web services with large candidate pools typically employ multi-stage recommender systems [13, 14, 7] (also known as cascade ranking [15, 16] process) to balance accuracy and efficiency under strict latency constraints. Each stage progressively narrows down the candidate set while employing increasingly sophisticated models. While most existing work focuses on independently optimizing individual stages (e.g., retrieval [17, 18] or reranking stage [7, 6]), recent studies explore joint optimization across the entire pipeline [16, 13, 19]. These approaches leverage later-stage models to guide earlier-stage learning, improving overall coherence and performance [20, 21]. However, our work takes a complementary direction—–using earlier-stage scores to regulate later-stage optimization.
@@ 2.2 Reranking Strategies
In sequential recommendation, re-ranking [6, 22, 8] aims to refine the initial ranking list generated by a previous retrieval stage. Existing re-ranking approaches can be broadly categorized into the following paradigms. Single-point re-scoring methods [23, 24] independently predict a refined score for each item and re-rank them accordingly. While efficient, they ignore dependencies between items, potentially leading to suboptimal list-level performance. Unlike single-point approaches, list-refinement methods [6, 8, 25] explicitly model mutual influences between items by treating the pre-ranked candidate list as input. Techniques such as Transformer-based architectures [6] capture user preference to optimize the list holistically. Generator-evaluator methods [22, 9] first generate multiple candidate lists and then evaluate them using a learned utility function to select the best one. They can effective generate a high-quality recommendation list but suffer from the increasing computational cost due to the generation-evaluation loop. Emerging techniques leverage diffusion models [11, 26] to generate refined item lists by iteratively denoising a ranking distribution, which excel in capturing complex user preferences.
@@ 2.3 Adversarial Learning
Adversarial learning [27] has emerged as a powerful technique for data augmentation in recommendation systems, improving the models robustness and generalization. By leveraging generative adversarial networks (GANs) [18] or adversarial training [28, 29], these methods synthesize high-quality user-item interactions or perturb existing data to enhance model robustness. For instance, IRGAN [18] employs a minimax game between a generator (sampling hard negative items) and a discriminator (distinguishing real vs. synthetic interactions) to improve ranking performance. Similarly, AdvIR [30] introduces adversarial perturbations to embedding spaces, forcing the model to learn more generalizable representations. These methods demonstrate that adversarial learning can not only augment training data but also enhance models’ robustness. Our method is designed to add noise by adversarial learning, which can enhance the denoise models (rerankers)’ robustness to uncertain noise in the pre-ranking scores.
@@ 3.1 Problem Formulation
For a given user request $\mathbf{u}$ (which may contain information about the candidate items $\mathcal{I}(u)$, user profile features, user interaction history), we observe the following during data collection: The candidate set $\mathcal{I}(u)=\{i_{1},i_{2},\cdots,i_{n}\}$ retrieved by the retriever; The corresponding (continuous) retriever scores of $\mathcal{I}(u)$ are $\mathbf{x}_{u}=[x_{i_{1}},x_{i_{2}},\cdots,x_{i_{n}}]\in[0,1]^{n}$, which essentially describes the use feedback likelihood in the view of the retriever; The ground truth user feedback (i.e., binary labels such as click, watch, share, etc.) of the candidate set, i.e., $\mathbf{z}_{u}=[z_{i_{1}},z_{i_{2}},\cdots,z_{i_{n}}]\in\{0,1\}^{n}$. Note that for recommender systems that further select K items after the reranking process, we may not observe the user feedback for the remaining items. Yet, from a probabilistic viewpoint, we can assume zeros as defaults since the item passing the reranking process is the premise of observing the user feedback signals.
The Noise Reduction Task: In our formulation, we denote the initial retriever as the prior $p_{\text{x}}(\mathbf{x}|\mathbf{u})$ and the learnable reranker $q_{\theta}(\mathbf{z}|\mathbf{x},\mathbf{u})$ as the conditional user feedback likelihood estimator . The $q_{\theta}$ predicts the user feedback given the retriever scores and the user context, i.e., $\hat{\mathbf{z}}_{u}\sim q_{\theta}(\mathbf{z}|\mathbf{x}_{u},\mathbf{u})$, and the learning goal adopts the general data likelihood maximization objective:
$\max p(\mathbf{z}_{u}|\mathbf{u})$ (1)
where the ground-truth $p(\mathbf{z}_{u}|\mathbf{u})$ is determined by the retriever $p_{\text{x}}$, the reranker $q_{\theta}$, and the user.
Note: For simplicity of the presentation, we omit the user condition term $\mathbf{u}$ for the remainder of the paper (e.g. we will denote the reranker as $q_{\theta}(\mathbf{z}|\mathbf{x}_{u})$ and simplify the learning goal as $\max p(\mathbf{z}_{u})$), since we adopt an analysis framework on the user request level.
@@ 3.2 Analytical Limitation of Direct Optimization Methods
which maximizes the user behavior alignment conditioned on the given retrieval scores. In practice, according to the binary nature of $\mathbf{z}_{u}$, we can simplify the learning objective as standard binary cross-entropy (BCE) loss. Theoretically, we can find that $\mathcal{L}_{\text{direct}}$ optimizes an upper bound of the negative data log likelihood (details in Appendix A.2):
@@
where $p_{\text{z|x}}$ represents the ground truth user feedback probability conditioned on retriever scores, and $p_{\text{x|z}}$ represents the ground truth posterior. Intuitively, smaller values of $L_{1}$ and $L_{2}$ make the solution of Eq.(2) more aligned with the goal of Eq.(1). However, these two terms are not fully controlled by the direct optimization framework and may generate a non-robust solution of $q_{\theta}$:
Irregular cases from retriever: The expectation term $L_{1}$ describes the difference between $q_{\theta}$ and the ground truth $p_{\text{z|x}}$ under the retriever prior. This term is non-positive and is minimized when $q_{\theta}(\mathbf{z}|\mathbf{x})=p(\mathbf{z}|\mathbf{x})$ for any variables. Though $L_{1}$ is expected to be small when the reranker’s predictions align with the user feedback in dominant cases, the retriever may happen to generate an $\mathbf{x}_{u}$ that the reranker fails to comprehend, inducing a significant divergence between the optimization of Eq.(2) and the learning goal of Eq.(1).
Regular cases with noisy retriever scores: The KL term $L_{2}$ describes the difference between the retriever score distribution $p_{\text{x}}$ and the obtained posterior given the knowledge of $\mathbf{z}_{u}$. This term is also non-positive and can be non-trivial due to the noisy nature of the retriever (as illustrated in Figure 1). This means that there might exist consistently significant misalignment between Eq.(2) and Eq.(1).
3.3 Implementing Noise Injection Posterior
To overcome the aforementioned limitations, we propose to complement the denoising reranker $q_{\theta}$ with a generative posterior $p_{\phi}(\mathbf{x}|\mathbf{z}_{u})$ that explicitly models the ground truth retriever score posterior distribution $p_{\text{x|z}}(\mathbf{x}|\mathbf{z}_{u})$. As described in Section 1, $p_{\text{x|z}}$ behaves as a noise-injecting model, so:
$\mathbf{x}_{u}^{\prime}=(1-\lambda_{c})\mathbf{z}_{u}+\lambda_{c}\bm{\epsilon}_{\phi},$ (4)
where the sampled error term $\bm{\epsilon}_{\phi}$ represents the noisy behavior of the posterior. To this end, we investigate two choices of noise samplers for $\bm{\epsilon}_{\phi}$:

### v2

# Denoising Neural Reranker for Recommender Systems

@@ Abstract
For multi-stage recommenders in industry, a user request would first trigger a simple and efficient retriever module that selects and ranks a list of relevant items, then the recommender calls a slower but more sophisticated reranking model that refines the item list exposure to the user. To consistently optimize the two-stage retrieval reranking framework, most efforts have focused on learning reranker-aware retrievers. In contrast, there has been limited work on how to achieve a retriever-aware reranker. In this work, we provide evidence that the retriever scores from the previous stage are informative signals that have been underexplored. Specifically, we first empirically show that the reranking task under the two-stage framework is naturally a noise reduction problem on the retriever scores, and theoretically show the limitations of naive utilization techniques of the retriever scores. Following this notion, we derive an adversarial framework DNR that associates the denoising reranker with a carefully designed noise generation module. The resulting DNR solution extends the conventional score error minimization loss with three augmented objectives, including: 1) a denoising objective that aims to denoise the noisy retriever scores to align with the user feedback; 2) an adversarial retriever score generation objective that improves the exploration in the retriever score space; and 3) a distribution regularization term that aims to align the distribution of generated noisy retriever scores with the real ones. We conduct extensive experiments on three public datasets and an industrial recommender system, together with analytical support, to validate the effectiveness of the proposed DNR.
@@ 1 Introduction
Recommender systems aim to personalize a list of items that maximizes user engagement signals, such as purchases in e-commerce, video plays in video platforms, and link sharing in social networks. To achieve a feasible solution with large-scale candidate pools and user requests, industrial solutions Covington et al. (2016); Ricci et al. (2022); Liu et al. (2025) typically employ a multi-stage architecture comprising a retriever and a reranker. The retriever, which may further decompose into several more sub-stages of retrieval and ranking modules, efficiently filters a large item pool down to a manageable candidate set of relevant items. Then the reranker adopts a more sophisticated but more accurate model that refines the candidate item set into an optimal ordering for the list-wise exposure. To optimize the two-stage framework towards a consistent goal of recommendation (e.g., aligning true user behavior or preference) across all stages, the majority of efforts have been focusing on learning a reranker-aware retriever Gallagher et al. (2019); Xu et al. (2024a). However, there is limited work on how to achieve a retriever-aware reranker.
@@
Formally, the goal of the latter reranking task Carbonell & Goldstein (1998); Jin et al. (2008); Pei et al. (2019); Liu et al. (2022) is to learn an ordering function over the retrieved candidates that maximizes user responses on the top-reranked items. Early studies propose to re-score the initial retrieval items Jin et al. (2008); Ai et al. (2018); Pei et al. (2019) by modeling the mutual influences among items in the candidate list. Recent state-of-the-art approaches find it more effective to formulate the reranking task as a list generation problem Shi et al. (2023); Ren et al. (2024); Lin et al. (2024), where the lists with better list-wise rewards are given higher generation probability. Yet, existing methods largely overlook the potential of initial retriever scores in the input space, which may offer rich prior information to enhance the reranking performance. In practice, a straightforward approach to leverage initial retriever scores is to include them as input features directly, which shows promise to improve reranking performance as in Figure 1(a). Still, as we will illustrate in section 3.2, this naive solution might be far from exploiting this retriever information effectively to align with user feedback, which is addressed by the following contributions in this work.
Formulating reranking as a noise reduction problem: The retriever stage typically employs a simpler model than the reranker to process large candidate pools, given computational budget constraints. In contrast, the reranker, operating on a much smaller candidate set, can use a more sophisticated architecture and thus achieves significantly higher accuracy (Figure 1(a), 1(b)). What follows is an empirical assumption that there is higher noise in the retriever scores than the reranker scores, which naturally indicates a noise reduction process of the reranking stage. This noise reduction pattern is more evident when we observe the changes in the embedding representations of ground truth target items. We show evidence by comparing Figure 1(c) and Figure 1(d), where the same set of retrieved candidates is not distinguished well by the retriever, but the target items are better predicted by the reranker, in terms of the embedding space. This further proves that the reranker behaves like a denoising model on the retriever’s noise. However, directly aligning the user behavior on the condition of the observed retriever score, ignoring the uncertainty of this noisy signal, may cause a misalignment with the user behavior alignment goal in terms of the entire system. Thus, we argue that learning a noise-aware reranker is crucial for achieving a retriever-aware reranker.
@@
The proposed solution: Motivated by the previous analysis, we propose an adversarial framework, which formulates a denoising neural reranker (DNR) with a carefully designed noise generation module, which introduces unseen noise to users’ feedback to augment the distribution modeling of retriever scores. We theoretically show that the user behavior alignment goal can be decomposed into three learning objectives in our solution framework: 1) a denoising objective that aims to denoise the noisy retriever scores under both observed and synthetic retriever scores; 2) an adversarial objective that encourages the noise generator to synthesize adversarial samples that are hard to denoise, which in turn improves the reranker’s effectiveness; and 3) a score distribution regularization term that aims to align the distribution between synthetic retriever scores with the real retriever scores.
We conduct empirical experiments on three public datasets and an industrial recommender system to verify the superiority of DNR against naive utilization strategies of retriever scores, as well as leading reranker models. Additionally, we find that the three proposed learning objectives in the adversarial denoising framework yield better recommendation performance than a single denoising objective.
@@ 2.1 Multi-stage Recommendation
Modern content-intensive web services with large candidate pools typically employ multi-stage recommender systems Qin et al. (2022); Zheng et al. (2024); Liu et al. (2022) (also known as cascade ranking Liu et al. (2017); Gallagher et al. (2019) process) to balance accuracy and efficiency under strict latency constraints. Each stage progressively narrows down the candidate set while employing increasingly sophisticated models. While most existing work focuses on independently optimizing individual stages (e.g., retrieval Wang et al. (2011; 2017) or reranking stage Liu et al. (2022); Pei et al. (2019)), recent studies explore joint optimization across the entire pipeline Gallagher et al. (2019); Qin et al. (2022); Xu et al. (2024b). These approaches leverage later-stage models to guide earlier-stage learning, improving overall coherence and performance Zhao et al. (2023); Xu et al. (2024a). However, our work takes a complementary direction—–using earlier-stage scores to regulate later-stage optimization.
@@ 2.2 Reranking Strategies
In sequential recommendation, re-ranking Pei et al. (2019); Huzhang et al. (2023); Ai et al. (2018) aims to refine the initial ranking list generated by a previous retrieval stage. Existing re-ranking approaches can be broadly categorized into the following paradigms. Single-point re-scoring methods Kang & McAuley (2018); Zhuang et al. (2018) independently predict a refined score for each item and re-rank them accordingly. While efficient, they ignore dependencies between items, potentially leading to suboptimal list-level performance. Unlike single-point approaches, list-refinement methods Pei et al. (2019); Ai et al. (2018); Xi et al. (2022) explicitly model mutual influences between items by treating the pre-ranked candidate list as input. Techniques such as Transformer-based architectures Pei et al. (2019) capture user preference to optimize the list holistically. Generator-evaluator methods Huzhang et al. (2023); Shi et al. (2023) first generate multiple candidate lists and then evaluate them using a learned utility function to select the best one. They can effective generate a high-quality recommendation list but suffer from the increasing computational cost due to the generation-evaluation loop. Emerging techniques leverage diffusion models Lin et al. (2024); Li et al. (2024) to generate refined item lists by iteratively denoising a ranking distribution, which excel in capturing complex user preferences.
@@ 2.3 Adversarial Learning
Adversarial learning Lowd & Meek (2005) has emerged as a powerful technique for data augmentation in recommendation systems, improving the models robustness and generalization. By leveraging generative adversarial networks (GANs) Wang et al. (2017) or adversarial training He et al. (2018); Tang et al. (2020), these methods synthesize high-quality user-item interactions or perturb existing data to enhance model robustness. For instance, IRGAN Wang et al. (2017) employs a minimax game between a generator (sampling hard negative items) and a discriminator (distinguishing real vs. synthetic interactions) to improve ranking performance. Similarly, AdvIR Park & Chang (2019) introduces adversarial perturbations to embedding spaces, forcing the model to learn more generalizable representations. These methods demonstrate that adversarial learning can not only augment training data but also enhance models’ robustness. Our method is designed to add noise by adversarial learning, which can enhance the denoising models (rerankers)’ effectiveness in denosing the pre-ranking scores to align with user feedback.
@@ 3.1 Problem Formulation
For a given user request $\mathbf{u}$ (which may contain information about the candidate items $\mathcal{I}(u)=\{i_{1},i_{2},\cdots,i_{n}\}$ retrieved by the retriever, user profile features, and user interaction history), we observe the following during data collection: 1) The corresponding (continuous) retriever scores of $\mathcal{I}(u)$ are $\mathbf{x}_{u}=[x_{i_{1}},x_{i_{2}},\cdots,x_{i_{n}}]\in[0,1]^{n}$, which essentially describe the user preference prediction in the view of the retriever; 2) The ground truth user feedback (i.e., binary labels such as click, watch, share, etc.) of the candidate set, i.e., $\mathbf{z}_{u}=[z_{i_{1}},z_{i_{2}},\cdots,z_{i_{n}}]\in\{0,1\}^{n}$. Note that for recommender systems that further select K items after the reranking process, we may never observe the positive user feedback for the remaining nKn-K items. From a probabilistic viewpoint, we can assume zero labels as defaults for these items, since passing the reranking process is the premise of observing the user feedback signals.
The Noise Reduction Task: In our formulation, the overall recommendation-feedback process $\mathbf{u}\rightarrow\mathbf{x}\rightarrow\mathbf{z}$ (where $\mathbf{x}$ and $\mathbf{z}$ are random variables for retriever scores and user feedback) involves an initial retriever $p_{\text{x}}(\mathbf{x}|\mathbf{u})$, which serves as the retriever score generation prior (and is assumed non-optimizable in the reranking task). The learnable reranker $q_{\theta}(\mathbf{z}|\mathbf{x},\mathbf{u})$ serves as the conditional user feedback likelihood estimator that predicts the user feedback $\hat{\mathbf{z}}$ given the retriever scores $\mathbf{x}$ and the user context $\mathbf{u}$, i.e., $\hat{\mathbf{z}}\sim q_{\theta}(\cdot|\mathbf{x}_{u},\mathbf{u})$. Given an observed feedback $\mathbf{z}_{u}$, the learning goal adopts the general data log likelihood maximization objective:
$\max_{\theta}\log p(\mathbf{z}_{u}|\mathbf{u},\theta)$ (1)
where the predicted likelihood $p(\mathbf{z}_{u}|\mathbf{u},\theta)$ for the observed user feedback $\mathbf{z}_{u}$ is determined by the retriever $p_{\text{x}}$, the reranker $q_{\theta}$, and the user request $\mathbf{u}$.
Note: For simplicity of the presentation, we omit the user condition term $\mathbf{u}$ for the remainder of the paper (e.g., we will denote the retriever as $p_{\text{x}}(\mathbf{x})$, the reranker as $q_{\theta}(\mathbf{z}|\mathbf{x}_{u})$, and simplify the learning goal as $\max_{\theta}\log p(\mathbf{z}_{u})$), since we adopt an analysis framework on the user request level.
@@ 3.2 Analytical Limitation of Direct Optimization Methods
which maximizes the user behavior alignment conditioned on the given retrieval scores. As a special case, according to the binary nature of $\mathbf{z}_{u}$, we can simplify the learning objective as standard binary cross-entropy (BCE) loss. Theoretically, we can find that $\mathcal{L}_{\text{direct}}$ optimizes an upper bound of the negative data log likelihood (details in Appendix B.2):
@@
where $p_{\text{z|x}}$ represents the ground-truth user feedback probability and $p_{\text{x|z}}$ represents the posterior distribution of retriever scores conditioned on user feedback, both terms are not related to the reranker $q_{\theta}$ and they are merely determined by the retriever prior and the user. Intuitively, smaller values of $L_{1}$ and $L_{2}$ make the solution of Eq.(2) more aligned with the goal of Eq.(1). However, these two terms are not fully controlled by the direct optimization framework, potentially causing significant misalignment between Eq.(2) and Eq.(1). This will lead to a discrepancy between the final reranking results and user feedback, and thus, the suboptimal recommendation performance.
3.3 Implementing Noise Generator for the Posterior
To overcome the aforementioned limitations, we propose to complement the denoising reranker $q_{\theta}$ with a noise generator $f_{\phi}(\cdot|\mathbf{z}_{u})$. Formally, we generate noise with $\bm{\epsilon}\sim f_{\phi}$ and add it to the user feedback to form the synthetic retriever score posterior $p_{\phi}(\mathbf{x}|\mathbf{z}_{u})$, taking insight from the reparameterization trick Kingma & Welling (2014):
$\mathbf{x}_{u}^{\prime}=(1-\lambda_{c})\mathbf{z}_{u}+\lambda_{c}\bm{\epsilon},$ (4)
where the generated noise $\bm{\epsilon}$ represents the noisy behavior in retriever scores $\mathbf{x}_{u}^{\prime}$, $\lambda_{c}$ represents the proportion of noise in the generated retriever scores. We denote the generation process as $\mathbf{x}_{u}^{\prime}\sim p_{\phi}$, which is analogous to the sampling process with ground truth posterior (assumed unknown and intractable), i.e., $\mathbf{x}\sim p_{\text{x|z}}$. For implementation, we investigate two choices of noise generator $f_{\phi}$: